# Optimal design of unimorph-type cantilevered piezoelectric energy harvesters using level set-based topology optimization by considering manufacturability

### v1

@@ 2.2 Formulation of piezoelectric phenomena
By weighting the modes by $\bar{\omega}_{i}$, we emphasized the lower-order modes.
In this study, we formulated an optimization problem that simultaneously minimizes distinct objective functions, that is, $F_{sb}$ for the substrate material domain $\Omega_{sb}$ and $F_{pe}$ for the piezoelectric material domain $\Omega_{pe}$, while maintaining a shared state field. This formulation incorporates the weighting factors $\alpha_{sb}$ and $\alpha_{pe}$ as demonstrated below Kim and Lee (2014):
$\begin{split}\underset{\Omega_{sb}}{\text{inf}}\qquad&F_{sb}=\alpha_{sb}F_{k}+(1-\alpha_{sb})F_{\omega}\\ \underset{\Omega_{pe}}{\text{inf}}\qquad&F_{pe}=\alpha_{pe}F_{k}+(1-\alpha_{pe})F_{\omega}\\ \text{subject}\;\text{to}\qquad&\text{governing}\;\text{equations}\;\text{for}\;\text{an}\;\text{open}\;\text{circuit},\\ \qquad&\text{governing}\;\text{equations}\;\text{for}\;\text{a}\;\text{short}\;\text{circuit}.\end{split}$ (9)
@@ 2.2.1 Minimum output voltage constraint
$V_{E}=\frac{Q}{C_{p}}=\frac{\int_{\Omega_{pe}}\bm{n}_{z}\cdot\nabla\varphi\ \text{d}\Omega}{\int_{\Omega_{pe}}\ \text{d}\Omega/(\varepsilon_{z}L_{z}^{2})},$ (10)
where $C_{p}$ signifies the capacitance of the piezoelectric energy harvester, $Q$ denotes the electrical charge elicited by the piezoelectric material, $\bm{n}_{z}$ represents a unit vector in the polarization direction, $\varepsilon_{z}$ is the polarization direction component of the piezoelectric material’s dielectric constant tensor, and $L_{z}$ is the thickness of the piezoelectric domain. The constraint on the minimum output voltage is formulated as follows:
$V_{E}=\frac{\int_{\Omega_{pe}}\bm{n}_{z}\cdot\nabla\varphi\ \text{d}\Omega}{\int_{\Omega_{pe}}\ \text{d}\Omega/(\varepsilon_{z}L_{z}^{2})}\geq\bar{V}_{minV},$ (11)
$\int_{\Omega_{pe}}\ \text{d}\Omega\leq\frac{\varepsilon_{z}L_{z}^{2}\int_{\Omega_{pe}}\bm{n}_{z}\cdot\nabla\varphi\ \text{d}\Omega}{\bar{V}_{minV}},$ (12)
@@ 2.3 Concept of topology optimization
Topology optimization is a type of structural optimization method. Structural optimization is used to obtain a structure $\Omega$ that minimizes or maximizes an objective function. The objective function often includes physical properties, such as stiffness Bendsøe and Kikuchi (1988); Allaire (2001), as well as thermal Zhou et al. (2019); Yamada et al. (2011); Jing et al. (2015); Wu et al. (2019); Miki and Yamada (2021), fluid Feppon et al. (2020); Guan et al. (2023), electromagnetic Choi and Yoo (2009); Yamada et al. (2013), and acoustic properties Wadbro and Berggren (2006); Sigmund and Jensen (2003); Noguchi and Yamada (2022, 2021). Therefore, the optimal structure is obtained under the assumption that the objective function satisfies the governing equations that describe the physical phenomena. The governing equations are considered as constraints in the optimization problem, and the basic structural optimization problem can be formulated as follows:
$\begin{split}\underset{\Omega}{\text{inf}}\qquad&F(u,\Omega)=\int_{\Omega}f(u)\mathrm{d\Omega}\\ \text{subject to}\qquad&\text{the governing equation system},\\ &\text{constraint equations},\end{split}$ (13)
@@
Next, we considered the application of topology optimization to the structural optimization problem (13). We introduced a domain $\Omega_{D}\subset\mathbb{R}^{n}(n=2\;\text{or}\;3)$ where the structure can be placed. Here, the domain $\Omega_{D}$ is called a fixed design domain because it does not change during the optimization process. The fixed design domain consists of a domain filled with the structure (hereafter referred to as the material domain) and a domain not filled with the structure (hereafter referred to as the void domain), and these domains are expressed by the characteristic function $\chi$, which is defined as follows:
$\chi(\bm{x}):=\left\{\begin{array}[]{ll}1\qquad\mathrm{for}\quad\bm{x}\in\Omega\\ 0\qquad\mathrm{for}\quad\bm{x}\in\Omega_{D}\backslash\Omega,\\ \end{array}\right.$ (16)
@@
In topology optimization, the problem described in (13) is reformulated as a material-distribution problem. This allows for topological changes, such as an increase or decrease in the number of holes, during the optimization procedure.
However, topology optimization problems are commonly ill-posed Allaire (2001); therefore, the space of admissible design should incorporate relaxation or regularization techniques to render the problem well-posed. A typical method based on the relaxation of space of an admissible design is the homogenization method Bendsøe and Kikuchi (1988). In this study, we used a level set-based topology optimization method Yamada et al. (2010) to transform an ill-posed problem into a well-posed problem. In this method, the boundary surface of the material domain is represented by the isosurface of a scalar function called the level set function, and changes in the level set function represent changes in the shape of the material domain. The topology optimization problem is regularized by ensuring the proper smoothness of the level set function. This method is described in subsection 2.3.1.
@@ 2.3.1 Level set-based topology optimization
In the level set-based method, the scalar function $\phi(\bm{x})\in H^{1}(\Omega_{D})$, also called the level set function, illustrated in the following equation is introduced to represent the shape:
$\left\{\begin{array}[]{lll}-1\leq\phi(\bm{x})<0&\mathrm{for}\quad\bm{x}\in\Omega_{D}\backslash\Omega\\ \phi(\bm{x})=0&\mathrm{for}\quad\bm{x}\in\partial\Omega\\ 1\geq\phi(\bm{x})>0&\mathrm{for}\quad\bm{x}\in\Omega,\\ \end{array}\right.$ (21)
@@
In the level set-based method, the topology optimization problem is formulated using the characteristic function defined in (24). However, as previously discussed, optimization that employs characteristic functions as design variables constitutes an ill-posed problem. Consequently, the level set-based method offers regularization to this optimization challenge, which is formulated as follows:
$\begin{split}\underset{\phi}{\text{inf}}\qquad&F_{R}(u,\phi)=\int_{\Omega_{D}}f(u)\chi_{\phi}(\phi)\mathrm{d\Omega}+\int_{\Omega_{D}}\frac{1}{2}\tau|\nabla\phi|^{2}\ \text{d}\Omega\\ \text{subject to}\qquad&\text{the governing equation system},\\ &\text{constraint equations},\end{split}$ (25)
@@
Next, we describe a method for updating the level set function. Assuming that the level set function is a function of fictitious time $t$, it is updated by the reaction-diffusion equation as follows:
$\begin{split}\frac{\partial\phi(\bm{x},t)}{\partial t}&=-K(\phi)\left\{-\tilde{c}F^{\prime}-\tau\nabla^{2}\phi(\bm{x},t)\right\},\\ \tilde{c}:&=\frac{c\int_{\Omega_{D}}\ \text{d}\Omega}{\int_{\Omega_{D}}|F^{\prime}|\ \text{d}\Omega},\end{split}$ (26)
where $K\in\mathbb{R}_{+}$ is the proportionality coefficient, $c\in\mathbb{R}_{+}$ is the normalization coefficient, and $F^{\prime}$ is the design sensitivity. In this study, we set $K=1.0$ and $c=2.0$.
@@ 3 Formulation of constraints for manufacturability
In this section, we describe the formulation of prerequisites related to manufacturing processes. This study set forth specific manufacturing criteria for piezoelectric energy harvesters, emphasizing configurations amenable to etching methodologies, predominantly utilizing the deep reactive ion etching (DRIE) apparatus Murakami et al. (2013) Laermer (1996). To this end, constraints were introduced to ensure structural uniformity in the thickness direction for both the piezoelectric and substrate domains. Such constraints prevent the emergence of configurations with apertures oriented in oblique or orthogonal directions relative to the processing trajectory. In addition, we advocated the imposition of a substrate-dependent constraint on the piezoelectric domain, precluding the occurrence of piezoelectric materials in areas devoid of a substrate. This constraint ensures that the piezoelectric materials do not exist in areas where the substrate is absent. The intricacies of each constraint are described below.

### v2

@@ 2.2.1 Matrix and vector formulation of piezoelectric phenomena
By weighting the performance $k_{i}$ by the modes $\bar{\omega}_{i}$, we emphasized the lower-order modes.
For a unimorph cantilevered piezoelectric energy harvester design, we perform the simultaneous optimization of the two objective functions $F_{k}$ and $F_{\omega}$ using a weighting factor $\alpha$ as follows:
$\underset{\Omega}{\text{inf}}\quad F=\alpha F_{k}+(1-\alpha)F_{\omega}.$ (10)
This approach allows us to obtain an optimized design that achieves enhanced power generation efficiency among solutions with the desired eigenfrequency. Furthermore, to optimize both the substrate and piezoelectric material regions, we employ the following method.
We formulate an optimization problem that simultaneously minimizes distinct objective functions for different material domains of the energy harvester, specifically, $F_{sb}$ for the substrate material domain $\Omega_{sb}$ and $F_{pe}$ for the piezoelectric material domain $\Omega_{pe}$. The substrate material domain $\Omega_{sb}$ consists of silicon, while the piezoelectric material domain $\Omega_{pe}$ is composed of PZT. The eigenvectors $\bm{u}_{oci}$ and $\bm{u}_{sci}$, as well as the electric potential $\varphi_{oci}$, are continuous across the boundary between $\Omega_{sb}$ and $\Omega_{pe}$.
The governing equations for each domain under open-circuit and short-circuit conditions are as follows. In the piezoelectric material domain $\Omega_{pe}$, the governing equations under open-circuit conditions are:
$\bm{C}^{E}_{pe}\bm{s}(\bm{u}_{oci})-\omega_{oci}^{2}\rho_{pe}\bm{u}_{oci}+\bm{e}^{T}\bm{\nabla}\varphi_{oci}=0\quad\text{in}\;\Omega_{pe}$ (11)
$\bm{e}\bm{s}(\bm{u}_{oci})-\bm{\varepsilon}^{S}\bm{\nabla}\varphi_{oci}=0\quad\text{in}\;\Omega_{pe}$ (12)
and under short-circuit conditions it is:
$\bm{C}^{E}_{pe}\bm{s}(\bm{u}_{sci})-\omega_{sci}^{2}\rho_{pe}\bm{u}_{sci}=0\quad\text{in}\;\Omega_{pe}.$ (13)
Similarly, in the substrate material domain $\Omega_{sb}$, the governing equations under open-circuit conditions is:
$\bm{C}^{E}_{sb}\bm{s}(\bm{u}_{oci})-\omega_{oci}^{2}\rho_{sb}\bm{u}_{oci}=0\quad\text{in}\;\Omega_{sb},$ (14)
and under short-circuit conditions is:
$\bm{C}^{E}_{sb}\bm{s}(\bm{u}_{sci})-\omega_{sci}^{2}\rho_{sb}\bm{u}_{sci}=0\quad\text{in}\;\Omega_{sb}.$ (15)
Here, $\bm{C}^{E}_{pe}$ and $\bm{C}^{E}_{sb}$ denote the elastic matrices of the piezoelectric and substrate materials, respectively, while $\rho_{pe}$ and $\rho_{sb}$ represent the densities of the piezoelectric and substrate materials. The eigenvector fields $\bm{u}_{oci}$ and $\bm{u}_{sci}$ are shared between the piezoelectric and substrate domains under open-circuit and short-circuit conditions, respectively.
This optimization incorporates weighting factors $\alpha_{sb}$ and $\alpha_{pe}$ to balance the contributions of each domain to the overall objective. The formulation of the objective functions $F_{sb}$ and $F_{pe}$, and the incorporation of the weighting factors $\alpha_{sb}$ and $\alpha_{pe}$, are demonstrated below, following the methodology proposed by Kim and Lee [22]:
$\begin{array}[]{rll}\underset{\Omega_{sb}}{\text{inf}}&F_{sb}=\alpha_{sb}F_{k}+(1-\alpha_{sb})F_{\omega}&\\ \underset{\Omega_{pe}}{\text{inf}}&F_{pe}=\alpha_{pe}F_{k}+(1-\alpha_{pe})F_{\omega}&\\ \text{subject to}&{\bm{C}^{E}_{pe}\bm{s}(\bm{u}_{oci})-\omega_{oci}^{2}\rho_{pe}\bm{u}_{oci}+\bm{e}^{T}\bm{\nabla}\varphi_{oci}=0}&\quad{\text{in}\;\Omega_{pe}}\\ &{\bm{e}\bm{s}(\bm{u}_{oci})-\bm{\varepsilon}^{S}\bm{\nabla}\varphi_{oci}=0}&\quad{\text{in}\;\Omega_{pe}}\\ &{\bm{C}^{E}_{pe}\bm{s}(\bm{u}_{sci})-\omega_{sci}^{2}\rho_{pe}\bm{u}_{sci}=0}&\quad{\text{in}\;\Omega_{pe}}\\ &{\bm{C}^{E}_{sb}\bm{s}(\bm{u}_{oci})-\omega_{oci}^{2}\rho_{sb}\bm{u}_{oci}=0}&\quad{\text{in}\;\Omega_{sb}}\\ &{\bm{C}^{E}_{sb}\bm{s}(\bm{u}_{sci})-\omega_{sci}^{2}\rho_{sb}\bm{u}_{sci}=0}&\quad{\text{in}\;\Omega_{sb}}\end{array}$ (23)
@@ 2.2.2 Minimum output voltage constraint
$V_{E}=\frac{Q}{C_{p}}=\frac{\int_{\Omega_{pe}}\bm{n}_{z}\cdot\nabla\varphi_{n}\ \text{d}\Omega}{\int_{\Omega_{pe}}\ \text{d}\Omega/(\varepsilon_{z}L_{z}^{2})},$ (24)
where $C_{p}$ signifies the capacitance of the piezoelectric energy harvester, $Q$ denotes the electrical charge elicited by the piezoelectric material, $\bm{n}_{z}$ represents a unit vector in the polarization direction, $\varepsilon_{z}$ is the component of the piezoelectric material’s dielectric constant tensor in the polarization direction, and $L_{z}$ is the thickness of the piezoelectric domain. $\varphi_{n}$ denotes the electric potential calculated using the modal superposition method. Details of the modal superposition method are described in A.
The constraint on the minimum output voltage is formulated as follows:
$V_{E}=\frac{\int_{\Omega_{pe}}\bm{n}_{z}\cdot\nabla\varphi_{n}\ \text{d}\Omega}{\int_{\Omega_{pe}}\ \text{d}\Omega/(\varepsilon_{z}L_{z}^{2})}\geq\bar{V}_{minV},$ (25)
$\int_{\Omega_{pe}}\ \text{d}\Omega\leq\frac{\varepsilon_{z}L_{z}^{2}\int_{\Omega_{pe}}\bm{n}_{z}\cdot\nabla\varphi_{n}\ \text{d}\Omega}{\bar{V}_{minV}},$ (26)
@@ 2.3 Concept of topology optimization
Topology optimization is a type of structural optimization method [39, 40, 41]. Structural optimization is used to obtain a structure $\Omega$ that minimizes or maximizes an objective function. The objective function often includes physical properties, such as stiffness [11, 39], as well as thermal [42, 43, 44, 45, 46], fluid [47, 48], electromagnetic [49, 50], and acoustic properties [51, 52, 53, 54].
2.3.1 Formulation of topology optimization
Typically in topology optimization methods, the optimal structure is obtained under the assumption that the objective function satisfies the governing equations that describe the physical phenomena. The governing equations are considered as constraints in the optimization problem, and the basic structural optimization problem can be formulated as follows:
$\begin{split}\underset{\Omega}{\text{inf}}\qquad&F(u,\Omega)=\int_{\Omega}f(u)\mathrm{d\Omega}\\ \text{subject to}\qquad&\text{the governing equations of the physical domain},\\ &\text{additional constraint equations},\end{split}$ (27)
@@ 2.3.1 Formulation of topology optimization
Next, we consider the application of topology optimization to the structural optimization problem (27). We introduce a domain $D\subset\mathbb{R}^{n}(n=2\;\text{or}\;3)$ where the structure can be placed. Here, the domain $D$ is called a fixed design domain because it does not change during the optimization process. The fixed design domain is a domain filled with the structure (hereafter referred to as the material domain) and a domain not filled with the structure (hereafter referred to as the void domain), and these domains are expressed by the characteristic function $\chi$, which is defined as follows:
$\chi(\bm{x}):=\left\{\begin{array}[]{ll}1\qquad\mathrm{for}\quad\bm{x}\in\Omega\\ 0\qquad\mathrm{for}\quad\bm{x}\in D\backslash\Omega,\\ \end{array}\right.$ (30)
@@
In topology optimization, the problem described in (27) is reformulated as a material distribution problem. This allows for topological changes, such as an increase or decrease in the number of holes, during the optimization procedure.
However, topology optimization problems are commonly ill-posed [39]; therefore, the space of admissible design should incorporate relaxation or regularization techniques to render the problem well-posed. A typical method based on the relaxation in space of the admissible designs is the homogenization method [11]. In this study, we use a level set-based topology optimization method [32] to transform an ill-posed problem into a well-posed problem. In this method, the boundary surface of the material domain is represented by the isosurface of a scalar function called the level set function, and changes in the level set function represent changes in the shape of the material domain. The topology optimization problem is regularized by ensuring the proper smoothness of the level set function. This method is described in subsection 2.3.2.
@@ 2.3.2 Level set-based topology optimization
In the level set-based method, the scalar function $\phi(\bm{x})\in H^{1}(D)$, also called the level set function, illustrated in the following equation is introduced to represent the shape [55]:
$\left\{\begin{array}[]{lll}-1\leq\phi(\bm{x})<0&\mathrm{for}\quad\bm{x}\in D\backslash\Omega\\ \phi(\bm{x})=0&\mathrm{for}\quad\bm{x}\in\partial\Omega\\ 1\geq\phi(\bm{x})>0&\mathrm{for}\quad\bm{x}\in\Omega,\\ \end{array}\right.$ (35)
@@
In the level set-based method, the topology optimization problem is formulated using the characteristic function defined in (38). However, as previously discussed, optimization that employs characteristic functions as design variables constitutes an ill-posed problem. Consequently, the level set-based method offers regularization to this optimization challenge, which is formulated as follows [56]:
$\begin{split}\underset{\phi}{\text{inf}}\qquad&F_{R}(u,\phi)=\int_{D}f(u)\chi_{\phi}(\phi)\mathrm{d\Omega}+\int_{D}\frac{1}{2}\tau|\nabla\phi|^{2}\ \text{d}\Omega\\ \text{subject to}\qquad&\text{the governing equations of the physical domain},\\ &\text{additional constraint equations},\end{split}$ (39)
@@
The optimization problem represented by Eq. (39), is reformulated using Lagrange’s method of undetermined multipliers. Let the Lagrangian be $L$ and the Lagrange multiplier of the governing equations be $v$, and that of the constraint equations be $\lambda$. The optimization problem is transformed into an unconstrained problem as follows:
$\underset{\phi}{\text{inf}}\quad L=\int_{D}f(u)\chi_{\phi}(\phi)\mathrm{d\Omega}+\int_{D}\frac{1}{2}\tau|\nabla\phi|^{2}\ \text{d}\Omega+v\;J(u)+\lambda\;G(u).$ (40)
Next, we describe a method for updating the level set function. Assuming that the level set function is a function of fictitious time $t$, it is updated by the reaction-diffusion equation as follows [32]:
$\begin{split}\frac{\partial\phi(\bm{x},t)}{\partial t}&=-K(\phi)\left\{-\tilde{c}F^{\prime}-\tau\nabla^{2}\phi(\bm{x},t)\right\},\\ \tilde{c}:&=\frac{c\int_{D}\ \text{d}\Omega}{\int_{D}|F^{\prime}|\ \text{d}\Omega}\;\;,\\ {F^{\prime}:}&{=\frac{\text{d}L}{\text{d}\phi}\;\;,}\end{split}$ (41)
where $K(\phi)\in\mathbb{R}_{+}$ is the proportionality coefficient, $c\in\mathbb{R}_{+}$ is the normalization coefficient, $L$ is the Lagrangian function, and $F^{\prime}$ is the design sensitivity defined as gradient of the Lagrangian function. $K(\phi)$ behaves as a coefficient to adjust the magnitude of sensitivity. In this study, we set $K=1.0$ and $c=2.0$.
@@ 3 Formulation of constraints for manufacturability
In this section, we describe the formulation of prerequisites related to manufacturing processes. This study sets forth specific manufacturing criteria for piezoelectric energy harvesters, emphasizing configurations amenable to etching methodologies, predominantly utilizing the deep reactive ion etching (DRIE) apparatus [9, 57]. Etching is a crucial step in the fabrication of microstructures, allowing for the precise removal of material. There are various etching techniques, primarily categorized into wet etching and dry etching. Wet etching involves chemical solutions to remove material, while dry etching uses plasma or gases. For the purposes of this study, we utilize the deep reactive ion etching (DRIE) method. Figure 2 outlines the DRIE process. DRIE is a highly anisotropic etching process that allows for the creation of complex structures in the planar direction using photolithography. This anisotropy means that etching can be precisely controlled in one direction, typically perpendicular to the substrate, while controlling etching in other directions is significantly more challenging. As a result, ensuring structural uniformity in the thickness direction is critical when designing structures using DRIE.
Figure 2: Steps of the DRIE process creating precise microstructures. (a) Application of resist: A layer of photoresist is applied to the surface of the material. (b) Photolithography: (1) A photomask is aligned over the photoresist-coated material. (2) UV exposure through the photomask transfers the pattern onto the photoresist. (3) The photoresist is developed, revealing the pattern. (c) Etching: (1) Etching begins in the exposed areas by exposing the surface to reactive ions. (2) Etching progresses to achieve the desired depth and pattern. (3) Etching is completed with precise control in the vertical direction. (d) Removal of resist: The remaining photoresist is removed, leaving behind the etched pattern.
To this end, constraints are introduced to ensure structural uniformity in the thickness direction for both the piezoelectric and substrate domains. Such constraints prevent the emergence of configurations with apertures oriented in oblique or orthogonal directions relative to the processing trajectory.
When etching is performed from only one side, the manufacturing constraints can be satisfied by imposing a constraint that the entire device has a uniform cross-section using a single level set function. This approach is mainly used for comparison in Section 5.1. On the other hand, when etching is performed from both sides, the structure does not need to have a uniform cross-section collectively. It is permissible for the piezoelectric material and the substrate to each have uniform cross-sections individually while maintaining overall manufacturability. By using the extended level set method [35], a type of multi-material topology optimization method, we can satisfy the constraints for etching from both sides while ensuring design flexibility.
In addition to ensuring structural uniformity, it is crucial to impose a substrate-dependent constraint on the piezoelectric domain. This is to preclude the occurrence of piezoelectric materials in areas devoid of a substrate, which can result in mechanical instability and reduced device performance. Without this constraint, the piezoelectric material may be deposited in areas where it lacks adequate support from the substrate, leading to potential delamination or cracking during the fabrication process. Figure 3 illustrates the difference between constrained and not constrained structures. Ensuring that the piezoelectric materials are only present where there is a supporting substrate helps maintain the mechanical integrity and reliability of the device.
In this section, we present different methods to model and address the etching fabrication constraint. Section 3.1 introduces methods to ensure that the cross-sectional shape of the device is uniform. Specifically, Section 3.1.1 describes the formulation of constraints for etching from a single side using the single level set function method, which is mainly used for comparative purposes. Section 3.1.2 presents the formulation of constraints for etching from both sides using the extended level set function method. Section 3.2 details the formulation of substrate-dependent constraints. Finally, Section 3.3 formulates the optimization problem incorporating both substrate-dependent constraints and the extended level set method.
Figure 3: Illustration of the substrate-dependent constraint on the piezoelectric domain. The left side shows a ”Constrained structure,” where the piezoelectric material is properly supported by the substrate, ensuring mechanical stability and integrity. The right side shows a ”Not constrained structure,” where piezoelectric material is deposited without substrate support, leading to potential delamination or cracking.
3.1 Constraint: the same cross-sectional shape
To ensure manufacturability via etching, it is crucial to impose constraints that maintain structural uniformity in the thickness direction. This section describes methods to achieve a uniform cross-sectional shape for the piezoelectric and substrate domains. The concept involves increasing the value of the regularization coefficient $\tau$ in the thickness direction to achieve uniform cross-sections. In this study, as shown in Figure 4, we have two different design domains: $D_{sb}$ for the substrate material and $D_{pe}$ for the piezoelectric material. The entire design domain is defined as $D=D_{pe}\cup D_{sb}$. We address two approaches to enforce this constraint: one using a single level set function for etching from a single side, and the other using the extended level set method for etching from both sides. Figure 5 illustrates the conceptual structures that satisfy each constraint. Structures as shown in Step (b) can be achieved by single-side etching, while structures as shown in Step (c) can be obtained through double-side etching.
Figure 4: The design domains used in this study. 3D view: Visualization of the 3D structure showing the division into the piezoelectric design domain ($D_{pe}$) and the substrate design domain ($D_{sb}$). Side view: Cross-sectional side view illustrating the components contained within each domain. $D_{pe}$ is the design domain for the piezoelectric material, while $D_{sb}$ is design domain for the substrate material.
Figure 5: Schematic illustration of the etching process for a unimorph cantilevered piezoelectric energy harvester. Initial configuration: Before any etching process begins. Step (a) Weight formation: The structure after etching from the backside to form the weight. Step (b) Beam and PZT formation: The structure after etching from the backside to form the silicon beam and PZT shapes. This step represents single-side etching. Step (c) Final PZT formation: The structure after etching from the front side to form the PZT shape differing from the beam. This step represents double-side etching.